OpenStudy (anonymous):

the quantities S and T are positive and are related by the equation S=k/T, where k is a constant. If the value of S increases by 50%, then the value of T decreases by what percent?

OpenStudy (anonymous):

50%

OpenStudy (anonymous):

50%

OpenStudy (anonymous):

let s= 10, t= 100 and consequently k = t*s = 1000 S increased by 50% to 10*1.5 = 15 15 = 1000/t t= 1000/15= 66.67 % decrease = 66.67- 100/100 = 33.33%